OpenStudy (anonymous):

The question is Factor 2x^2 = 3x - 5. I have no idea what to do...

OpenStudy (anonymous):

Hello, Welcome to Open Study!!!!!

OpenStudy (anonymous):

Let 2x^2=3x-5

OpenStudy (anonymous):

2x^2-3x+5=0

OpenStudy (anonymous):

(-)(-).....

OpenStudy (anonymous):

(2x-5)(x-1)=0

OpenStudy (anonymous):

hence x =5/2, and 1!!!!!

OpenStudy (anonymous):

Happy Holidays!!!!!